# Author guidelines for ISBI proceedings manuscripts

### 15 Dec 2023 v1

@@ 1 Introduction
The electrocardiogram (ECG) is a crucial time series data that provides valuable insights into the detection and diagnosis of cardiac diseases. The morphology observed in each lead of ECG provides insight into the electrical activity occurring within specific segments of the heart. This information proves invaluable in identifying abnormal myocardial conditions [1]. However, focusing solely on the classification of specific diseases through ECG analysis may result in the failure to detect various types of abnormal signals that were not encountered during training. Therefore, our attention shifts towards the detection of anomalies in ECG signals, a strategy with the potential to identify abnormal cardiac conditions in patients, regardless of the diversity and rarity of these conditions.
@@ 4 Experiments
We compare our TSRNet with several SOTA methods in time series anomaly detection, e.g. DAGMM [2], MAD-GAN [3], USAD [4], TranAD [9], Anomaly Transformer [5], Zheng et al. [17], BeatGAN [10], and Jiang et al. [11]. All the performance results are excerpted from Jiang et al. [11]. Those SOTA methods belong to the unsupervised time series anomaly detection approach. As a result, we regard them as the most suitable options for comparison in order to highlight the effectiveness of our proposed method.
@@
Table 3: Ablation study on the effectiveness of each individual modality (left) and the effectiveness of the proposed Peak-based Error (right).
@@
Moving to Table 3 (right), we assess the efficacy of our proposed Peak-based Error during the inference process. The findings indicate that the inclusion of the Peak-based Error is instrumental in achieving high performance.

### 05 Mar 2024 v2

@@ 1 Introduction
The electrocardiogram (ECG) is a crucial time series data that provides valuable insights into the detection and diagnosis of cardiac diseases. The morphology observed in each lead of ECG provides insight into the electrical activity occurring within specific segments of the heart. This information proves invaluable in identifying abnormal myocardial conditions [1, 2, 3, 4]. However, focusing solely on the classification of specific diseases through ECG analysis may result in the failure to detect various types of abnormal signals that were not encountered during training. Therefore, our attention shifts towards the detection of anomalies in ECG signals, a strategy with the potential to identify abnormal cardiac conditions in patients, regardless of the diversity and rarity of these conditions.
@@ 4 Experiments
We compare our TSRNet with several SOTA methods in time series anomaly detection, e.g. DAGMM [5], MAD-GAN [6], USAD [7], TranAD [12], Anomaly Transformer [8], Zheng et al. [18], BeatGAN [13], and Jiang et al. [14]. All the performance results are excerpted from Jiang et al. [14]. Those SOTA methods belong to the unsupervised time series anomaly detection approach with the same setting as our method. As a result, we regard them as the most suitable options for comparison in order to highlight the effectiveness of our proposed method.
@@
Table 3: Ablation study on the effectiveness of each modality (left) and the effectiveness of the proposed Peak-based Error (right).
@@
Moving to Table 3 (right), we assess the efficacy of our proposed Peak-based Error during the inference process. We compare the result when inferring with Peak-based Error to the result when inferring the original objective function [14]. The findings indicate that the inclusion of the Peak-based Error is instrumental in achieving high performance.
@@ 5 Conclusion
6 Acknowledgement
Nhat-Tan Bui, Thinh Phan, and Ngan Le are supported by the National Science Foundation (NSF) under Award No OIA-1946391 RII Track-1, NSF 1920920 RII Track 2 FEC, NSF 2223793 EFRI BRAID, NSF 2119691 AI SUSTEIN, NSF 2236302. Minh-Triet Tran is sponsored by Vietnam National University Ho Chi Minh City (VNU-HCM) under grant number DS2020-42-01. Dinh-Hieu Hoang is funded by Vingroup Joint Stock Company and supported by the Domestic Master/ PhD Scholarship Programme of Vingroup Innovation Foundation (VINIF), Vingroup Big Data Institute (VINBIGDATA), code VINIF.2022.ThS.JVN.04.